OpenStudy (anonymous):

Simplify: (2 - 3i)^2(2 + 3i)^2

OpenStudy (anonymous):

first rearrange as \[((2+3i)(2-3i))^2\]

OpenStudy (anonymous):

multiplying a number by its complex conjugate is easy. you get \[(a+bi)(a-bi)=a^2+b^2\]

OpenStudy (anonymous):

in this case you get \[(4+9)^2=13^2=169\]